OpenStudy (anonymous):

Math help! Lets say i have 28 soccer balls. 22 is red and the rest are pink. What is the percentage of red. I just need help trying to see how much "1" is in percentage if the whole is 28.

OpenStudy (tkhunny):

If you have only pink and red soccer balls, there is something wrong with you. 1 of 10 is 10% 1/10 = 0.10 = 10% 1 of 20 is 5% 1/20 = 0.05 = 5% 1 of 40 is 2.5% 1/40 = 0.025 = 2.5%

OpenStudy (anonymous):

just a metaphor ;)

OpenStudy (anonymous):

so should I use these as benchmarks?

OpenStudy (anonymous):

I still dont understand should i round>

OpenStudy (tkhunny):

Not really. You should see how to calculate a percentage. It's a division problem. 3 of 10 is 30% 3/10 = 0.30 = 30% 3 of 20 is 15% 3/20 = 0.15 = 15% 3 of 40 is 7.5% 3/40 = 0.075 = 7.5%

OpenStudy (tkhunny):

If rounding is the rule of the day, then go ahead. If not, then don't. 1/28 is what percent? 3/28 is what percent?

OpenStudy (tkhunny):

1/28 = 0.03571428571428571428571428571429 = 3.57% or so maybe 3.57143% 3/28 = 0.10714285714285714285714285714286 = 10.71% or so maybe 10.7143%